# Prompt Injection Vulnerability of Consensus Generating Applications in Digital Democracy

### v1

@@ 5 Results
This section presents our results on testing the vulnerability of LLMs to prompt injection attacks, and the effectiveness of DPO to produce LLM that are robust against this kind of attacks. All our experiments were run on external servers using Google Colab and the OpenAI Python library.
@@ 5.1 Vulnerability to Prompt-Injections
Figure 4: Effectiveness of prompt-injection attacks by taxonomy dimensions: ignore/completion, framing, and rethorical strategy.
Figure 5: Robustness to prompt-injection attacks via DPO by taxonomy dimensions: ignore/completion, framing, and rethorical strategy.
@@
Figures 4 (a), (b), (d), and (e) present a breakdown of the effectiveness of prompt-injection attacks, distinguishing between the ignore/completion attacks and framing (support/criticism) axes of our taxonomy. Each 3×3 confusion matrix captures how LLM-generated consensus’ verdicts shift when exposed to attacks, with separate matrices for support attacks and criticism attacks.
@@
We find that consensus generating in the tested LLM is more vulnerable to prompt-injections that are framed negatively (as criticism) and target ambiguous consensus. In terms of rhetorical strategy, explicit commands and superficially rational arguments (”Imperative Order” and ”Impossibility of Agreement”) outperform emotional language or fake statistics.
@@
Note that as injections are only applied when the participant’s verdict aligns with the framing of the attack—e.g., support injections are only applied to initially agreeing opinion texts—, the number of prompts tested for support attacks (2,448) and criticism attacks (2,568) differ.
@@
Figure 4 (a) shows the effects of ignore-support attacks: even with this simple injection, 50.4% of originally disagreeing and 65.3% of ambiguous consensus statements shift to agreement. This effect is stronger with completion-support attacks (Figure 4 (b), where 76% of originally disagreeing and 73.6% of ambiguous statements are steered toward agreement, suggesting completions are more effective than ignore attacks.
@@
Figures 4 (d) and (e) show results for criticism attacks using ignore and completion injections, respectively. Ignore injections produce the unintended effect of shifting ambiguos statements toward agreement. In fact, 27.2% of agreeing and 76.4% of ambiguous statements shift toward agreement, contrary to the attack’s intent—likely due to interactions with the LLM’s priors. In contrast, completion-criticism attacks are more effective, shifting 64.2% of agreeing and 95.8% of ambiguous statements to disagreement. With an ASR of 46.3%, they are the most successful rethorical strategy.
@@
Figures 4 (c) and (f) break down attack effectiveness by the rethorical strategy dimension of our taxonomy, revealing that rational-sounding and explicit instructions outperform emotional language and fabricated statistics. Interestingly, our results are consistent with the findings of Zeng et al., 2024. Among support attacks, the strategy ”Impossiblity of Agreement”—based on declaring that a consensus is too difficult to find—emerges as the most effective, with ASRs of 47.8% in ignore and 56.4% in completion prompt-injections. By contrast, the success of criticism attacks depends more on the ignore/completion dimension and less on the rethorical strategy.
@@
We replicated these experiments using GPT-4.1 Nano (see supplementary material). We still find that completion are more effective than ignore injections for support (32.15% vs. 27.73%) and criticism attacks (33.99% vs. 24.57%). Ambiguous statements remain fragile: less than 31.9% withstand such attacks under criticism attacks and less 23.6% under support. The success patterns observed in the rethorical strategy dimension are also replicated.
@@ 5.2 DPO as Protection Against Prompt Attacks
We reproduce the analyses from the previous section using consensus statements generated by the LLM aligned with DPO, both with and without prompt-injections. Results are shown in Figure 5, mirroring the structure of Figure 4.
Overall, alignment via DPO enhances our LLM’s robustness to prompt-injection attacks. The effects are stronger in cases where injection texts are based on rational-sounding language and explicit instructions. Unsurprisingly, ambiguous consensus statements remain the most easily shifted, showing that the alignment is not a sufficient protection strategy against prompt-injection attacks in these cases.
Note that, since injections are only applied when the participant’s verdict aligns with the framing of the attack—e.g., support injections are only applied to initially agreeing opinion texts—, the number of prompts tested for support attacks (2,784) and criticism attacks (3,408) differ.
DPO yields marginal robustness improvements for ignore prompt-injections. While alignment helps preserve the original verdicts in clear-cut cases—especially among disagreeing consensus inputs—our LLM struggles to maintain ambiguous verdicts. Under support attacks the ASR decreases from 25.65% to 19.54% (panel (a)) and for criticism attacks the ASR falls modestly from 28.66% to 21.19% (panel (d)).
In contrast, completion prompt-injections exhibit the strongest robustness gains under DPO alignment, particularly in the face of criticism attacks. Figure 5 (e) shows that the attack success rate (ASR) drops by more than half, from 46.34% to 19.37%. This improvement is driven by a sharp decline in verdict reversals: only 21.5% of originally agreeing consensus statements flip to disagreement, compared to 64.2% in the unaligned LLM. This is less true of ambiguous statements, where 45.8% still shift toward disagreement under attack—down from 95.8% without alignment.
Figures 5 (c) and (f) break down attack effectiveness by the rethorical strategy dimension. We find that DPO substantially eliminates the asymmetries in ASRs across the taxonomy dimensions. After alignment, most strategies converge toward ASR values between 16.4% and 20.3%.
These findings replicate with GPT-4.1 Nano (see supplementary material), although with lower success. After DPO, the overall ASRs drop from values between 24.57% and 34% to values between 26.23% and 28.29% across all ignore/completions and framings.
@@ 5.3 Extended Analysis
To test the generality of our findings, we replicate our analysis using 26,502 candidate statements submitted to the Smartvote voting advice application during Switzerland’s national parliamentary elections, addressing 374 public policy questions (Stammbach et al. 2024). For each combination of political party (26), language (French, German, Italian), and public-policy question (374), we simulate consensus statements by aggregating 5-40 opinion texts, resulting in 96,384 test set prompts. We find similar ASR values—between 19.85% and 26.03%—, and that effectiveness of single attacks scales to cases with >30>30 participants. Finally, results using (i) Syntatic Dependency Analysis (with no LLMs), and (ii) Deliberative Alignment—based on adding security policies to reasoning before generating statements (Guan et al. 2024)—as benchmarks show no better results than DPO (see supplementary material).
@@ 6 Conclusion and Future Work
Large language models (LLMs) are increasingly employed to generate consensus statements in support of group deliberation, representing one of the most promising application of AI in digital democratic systems (Konya et al. 2025; Small et al. 2023; Tessler et al. 2024). In this paper we systematically explored the space of prompt-injection attacks in this setting, introducing a taxonomy of injections that enables fine-grained analysis of both attack efficacy and mitigation strategies.
Our analysis reveals fundamental vulnerabilities, particularly when injections are framed negatively (disagreement) and when the target is ambiguous consensus. We also show that completion strategies—which steer the LLM toward a specific text—are more effective than attempts to suppress opposing views. These attack types are especially concerning because they resemble inputs any participant might provide, without employing complex injection techniques or even unintentionally. While alignment techniques improve robustness, they do not fully address the issue, underscoring the need for further research on secure and reliable deliberation with LLMs.
@@ 7 Acknowledgments
Appendix A Rethorical Strategies
@@ Appendix B BERT Classifier
We used GPT-4o to classify 1,690 opinion texts into one of three verdicts regarding public-policy questions: ”Agreement”, ”Disagreement”, or ”Ambiguous”. Using these tags, we fine-tuned a BERT model to replicate the labeling process at a much lower cost, achieving an F1 score and accuracy of 98% on a set of 600 opinion texts not seen during fine-tuning. After several experiments, the highest accuracy was reached with a learning rate of 2e-5, a batch size of 16 observations, 5 epochs and a weight decay of 0.
@@ Appendix C Performance of default LLMs
Using our BERT-based classifier, we evaluated each consensus statement generated by different LLMs and measured the extent to which the verdict of each statement aligned with the majority opinion of participants for each question.
@@ Appendix D Direct Policy Optimization (DPO)
E Quality Metrics: ROUGE-L F1 Scores
E Quality Metrics: BERTScore F1
F Semantic Diversity Metrics: Jaccard Similarity
G Vulnerability and Robustness with GPT 4.1 Nano
(a) Vulnerability to prompt-injection attacks by taxonomy dimensions: ignore/completion, framing, and rhetorical strategy.
(b) Robustness against prompt-injection attacks via DPO by taxonomy dimensions: ignore/completion, framing, and rhetorical strategy.
Figure 11: GPT 4.1 Nano vulnerability and robustness analysis across attack taxonomy dimensions
H Results with Smartvote
We validate our findings using a dataset of comments submitted to Smartvote, a popular Swiss voting advice application, to simulate intra-party consensus statements. From a database containing approximately 100,000 comments written by candidates running for national parliament in Switzerland (Stammbach et al. 2024), we utilize 26,502 comments to examine LLMs’ vulnerabilities in consensus formation scenarios involving 5-40 opinion texts.
We define consensus tasks for each unique combination of political party (26 parties), language (French, German, Italian), and public-policy question (374 questions). This framework allows us to analyze the effects of the same taxonomy dimensions established in the main text while examining how ASR values vary with the number of participants involved in the consensus formation process. A multilingual BERT model was fine-tuned for defining the verdicts (Boizard et al. 2025) with default hyperparameters. After filtering out prompts that did not satisfy minimum requirements, 24,216 were used for analyzing support and 23,976 for criticism attacks.
(a) Vulnerability to prompt-injection attacks by taxonomy dimensions, SmartVote dataset.
(b) ASR as a function of the number of participants.
Figure 12: Smartvote dataset analysis: vulnerability patterns and participant scaling effects
I Benchmarks
I.1 Syntactic Dependency Analysis
@@
Using this procedure, we find that 53.46% of injection texts containing Emotional Appeals, 1.53% containing False Authority injections, 10.76% containing imperative orders, 7.69% containing ”Impossibility of Agreement” injections, and 8.65% containing injections based on False Statistics. Additionally, 43.73% of benign texts written by participants in the dataset from (Tessler et al. 2024) also contain imperatives, revealing a key weakness of this approach in distinguishing rhetorical and persuasive content.

### v2

@@ 5.1 Vulnerability to Prompt-Injections
Figures 4 (a), (b), (d), and (e) present a breakdown of the effectiveness of prompt-injection attacks along the axes of our taxonomy for LLaMA 3.1 8B Instruct (see Appendix F for results with GPT 4.1 Nano and Apertus 8B). Each 3×3 confusion matrix captures how LLM-generated consensus’ valences shift when exposed to attacks, with separate matrices for support attacks and criticism attacks.
@@
We find that consensus generation in the tested LLM is more vulnerable to prompt-injections that are framed negatively (as criticism) and target ambiguous consensus. In terms of rhetorical strategy, explicit commands and superficially rational arguments (”Imperative Order” and ”Impossibility of Agreement”) outperform emotional language or fake statistics.
@@
Note that as injections are only applied when the participant’s valences aligns with the framing of the attack—e.g., support injections are only applied to initially agreeing opinion texts—, the number of prompts tested for support attacks (2,448) and criticism attacks (2,568) differ.
@@
Figure 4 (a) shows the effects of ignore-support attacks: even with this simple injection, 40.8% of originally disagreeing and 75% of ambiguous consensus statements shift to agreement. This effect is stronger with completion-support attacks (Figure 4 (b), where 59.3% of originally disagreeing and 86.1% of ambiguous statements are steered toward agreement, suggesting completions are more effective than ignore attacks.
@@
Figures 4 (d) and (e) show results for criticism attacks using ignore and completion injections, respectively. Ignore injections produce the unintended effect of shifting ambiguous statements toward agreement. In fact, 15.7% of disagreeing and 69.4% of ambiguous statements shift toward agreement, contrary to the attack’s intent—likely due to interactions with the LLM’s priors. In contrast, completion-criticism attacks are more effective, shifting 75.1% of agreeing and 97.2% of ambiguous statements to disagreement. With an ASR of 45.47%, they are the most successful rhetorical strategy.
@@
Figures 4 (c) and (f) break down attack effectiveness by the rhetorical strategy dimension of our taxonomy, revealing that rational-sounding and explicit instructions outperform emotional language and fabricated statistics. Interestingly, our results are consistent with the findings of Zeng et al., 2024. Among support attacks, the strategy ”Impossibility of Agreement”—based on declaring that a consensus is too difficult to find—emerges as the most effective, with ASRs of 34.6% in ignore and 54.1% in completion prompt-injections. By contrast, the success of criticism attacks depends more on the ignore/completion dimension and less on the rhetorical strategy.
@@
We replicated these experiments using GPT-4.1 Nano and Apertus 8B (see Appendix F). We still find that completion are more effective than ignore injections for support (32.15% vs. 27.74% in GPT 4.1; 23.86% vs. 19.04% in Apertus) and criticism attacks (33.98% vs. 24.58% in GPT 4.1; 42% vs. 22.37% in Apertus). Ambiguous statements remain fragile: less than 31.9% withstand such attacks under criticism attacks and less 23.6% under support. The success patterns observed in the rhetorical strategy dimension are also replicated.
In Appendix G we present results for evaluation metrics related to textual quality (ROUGE-L F1 score, F1 BERTScores) and in Appendix H for semantic diversity (Jaccard Similarity).
@@ 5.2 DPO as Protection Against Prompt Attacks
Figure 5 reports a comparative evaluation of robustness defenses against prompt-injection attacks across two LLMs—LLaMA 3.1 8B Instruct and GPT 4.1 Nano. Results are organized along two axes of our taxonomy: ignore/completion and the framing of the attack (support vs. criticism), with the ASR serving as the main performance indicator.
For LLaMA 3.1 8B Instruct (panel a), our simulations of consensus statements with prompt injection attacks reveal a LLM partially robust under alignment. DPO alone reduces ASRs by roughly seven to eight percentage points relative to the undefended baseline, with the strongest effect observed in completion-criticism attacks (from 45.5% to 16.8%). Still, less effective attacks in the undefended baseline prove more challenging, reaching baseline ASRs ranging from 16.3% to 19.5% in the rest of ignore/completion and framing cases. When layered defenses are introduced, different patterns emerge. DPO combined with SecAlign++ (MetaSecAlign) brings worse results than the undefended baseline for support framings but is more effective under criticism-completion attacks, lowering the ASR to 23.5%. The integration of DPO with defensive tokens and reformatting opinion texts (”DataFilter”) yields more consistent protection across settings, reducing average ASRs to values ranging from 15.6% to 20.7% for both support and criticism completions—an improvement of roughly 10 percentage points from the undefended baseline. Adding reinforcement learning with GRPO to defensive tokens did not yield better results, with ASRs ranging from 15.6% to 23.9%. These results suggest that token-level embeddings optimization and input reformatting offer a stabilizing effect when alignment alone proves insufficient.
For GPT 4.1 Nano (panel b), robustness follow a different pattern. The undefended baseline ASRs are relatively uniform, spanning from 24.6% to 34%. In this case, standard DPO alignment offers minimal gains, keeping ASRs between 26.3% and 28.3%. Yet, when DPO is coupled with chain-of-thought (CoT) reasoning—that is, following the Deliberative Alignment’s approach (Guan et al. 2024)—, the defenses strengthen dramatically: ASRs fall to between 6.9% and 11.5% across all attack categories. This reduction—equivalent to 60–80% relative to baseline—indicates that chains-of-thought based alignment mechanisms, by reinforcing the LLMs’ deliberative structure, provide a more durable form of robustness in larger, instruction-tuned architectures in DD settings. Adding input reformatting on top of CoT yields worse results for criticism attacks (19.5% and 17.5%), but since consensus statement valences are predicted as ambiguous (see DataFilter, Appendix E), this represents a desirable property.
Taken together, these results underscore that robustness is LLM-dependent when using DPO alignment. Smaller models, such as LLaMA 3.1 8B Instruct, benefit most from token-level embedding optimization, while larger models, like GPT-4.1 Nano, gain more from chain-of-thought augmented defenses. Input reformatting provides further improvements grounded in safe reinterpretations. One plausible explanation is that CoT enables LLMs to identify malicious instructions expressed through highly indirect language, while input reformatting removes linguistic attack cues. In contrast, other layered defenses primarily target explicit attacks—the same focus underlying benchmarks such as Qwen3Guard and GPT-OSS Guard, which exhibit poor protection (see Appendix E).
An over-refusal rate analysis (Appendix J), comparing consensus statement valences with and without alignment in the absence of prompt injection attacks, shows that defensive tokens provide the lowest over-refusal risk, with ASR values ranging between 12% and 16.1%.
@@ 5.3 Extended Analysis
To test the generality of our findings, we replicated the analysis on 26,502 candidate statements from the Smartvote platform during Switzerland’s parliamentary elections, covering 374 policy questions (Stammbach et al. 2024). For each combination of party, language, and policy question, we generated within-party consensus statements by aggregating 5–40 opinions, yielding 127,920 test prompts that varied in framing, ignore/completion, and rhetorical strategy. Results show ASRs remain nonzero and indicate that the effectiveness of single attacks scales to groups with more than 30 participants. Finally, replacing LLM-based benchmarks with Syntactic Dependency Analysis (SDA) as a lightweight detector produces even higher ASRs than DPO (see Appendix E.1), underscoring the need to evaluate robustness with state-of-the-art approaches.
Finally, we conduct additional analysis in Appendix K using a measure we call Directional Attack Success Rate (DirASR), defined as the proportion of consensus pairs in which a single prompt injection shifts the LLM’s valence away from the direction the attack is designed to promote. Specifically, DirASR measures the share of consensus pairs that fall in the top-left and top-center cells as a percentage of pairs in the left column for criticism attacks, and the share falling in the center-right and bottom-right cells as a percentage of pairs in the right column for support attacks, as shown in Figure 2. We find that DPO and input reformatting yield DirASR values below 10% across different LLMs, framings, and ignore/completion categories, while the remaining defenses yield lower effectiveness. Further exploration of other layered defenses is needed.
@@ 6 Conclusion and Future Work
LLMs are increasingly employed to generate consensus statements in support of group deliberation, representing one of the most promising application of AI in digital democratic systems (Konya et al. 2025; Small et al. 2023; Tessler et al. 2024). In this paper we systematically explored the space of prompt-injection attacks in this setting, introducing a simple taxonomy of injections that enables fine-grained analysis of both attack efficacy and mitigation strategies.
Our analysis reveals fundamental vulnerabilities, specially when injections are framed negatively (disagreement) and when the target is ambiguous consensus. We also show that completion strategies—attacking LLMs by recreating a false consensus statement–are more effective than deliberate attempts to suppress viewpoints. These attack types are especially concerning because they resemble inputs any participant might provide, without employing sophisticated injection techniques or technical expertise. While alignment techniques improve robustness, they do not fully address the issue, underscoring the need for further research on secure and reliable deliberation with LLMs.
@@ 7 Acknowledgments
Appendix A Rhetorical Strategies
@@ Appendix B BERT Classifier
We used GPT-4o to classify 1,690 opinion texts into one of three valences regarding public-policy questions: ”Agreement”, ”Disagreement”, or ”Ambiguous”. Using these tags, we fine-tuned a BERT model to replicate the labeling process at a much lower cost, achieving an F1 score and accuracy of 98% on a set of 600 opinion texts not seen during fine-tuning. After several experiments, the highest accuracy was reached with a learning rate of 2e-5, a batch size of 16 observations, 5 epochs and a weight decay of 0.
@@ Appendix C Performance of default LLMs
Using our BERT-based classifier, we evaluated each consensus statement generated by different LLMs and measured the extent to which the valence of each statement aligned with the majority opinion of participants for each question.
@@ Appendix D Direct Policy Optimization (DPO)
E Robustness Methodologies and Benchmarks
E.1 Syntactic Dependency Analysis
@@
Using this procedure, we identified only 53.46% of Emotional Appeals injections, 1.53% of False Authority injections, 10.76% of imperative orders, 7.69% of ”Impossibility of Agreement” injections, and 8.65% of False Statistics injections. Additionally, 43.73% of benign texts written by participants in the Tessler dataset were idenfitied as injections, revealing a key weakness of this approach in distinguishing rhetorical and persuasive content.
@@ E.2 Deliberative Alignment
F Vulnerability Results with GPT 4.1 Nano and Apertus 8B Instruct
(a) GPT 4.1 Nano
(b) Apertus 8B
Figure 10: GPT 4.1 Nano and Apertus 8B vulnerability analysis across attack taxonomy dimensions: ignore/completion, framing, and rhetorical strategy
G Quality Metrics: ROUGE-L F1 Scores
G Quality Metrics: BERTScore F1
H Semantic Diversity Metrics: Jaccard Similarity
I Results with Smartvote
We validate our findings using a dataset of comments submitted to Smartvote, a popular Swiss voting advice application, to simulate intra-party consensus statements. From a database containing approximately 100,000 comments written by candidates running for the national parliament in Switzerland (Stammbach et al. 2024), we utilize 26,502 comments to examine LLMs’ vulnerabilities in consensus formation scenarios involving 5-40 opinion texts.
We define consensus tasks for each unique combination of political party (26 parties), language (French, German, Italian), and public-policy question (374 questions). This framework allows us to analyze the effects of the same taxonomy dimensions established in the main text while examining how ASR values vary with the number of participants involved in the consensus formation process. We fine-tuned a multilingual BERT model (Boizard et al. 2025) with default hyperparameters for defining valences using 35,665 “Agree”/“Disagree” Swiss candidates ratings and “Ambiguous” ratings from GPT-4o. After filtering out prompts that did not satisfy minimum requirements, 24,216 were used for analyzing support and 23,976 for criticism attacks.
(a) ASR as a function of the number of participants using LLaMA 3.1 8B Instruct.
(b) ASR as a function of the number of participants using Apertus 8B Instruct.
Figure 14: Smartvote dataset analysis: participant scaling effects
J Overrefusal Rates
(a) LLaMA 3.1 8B Instruct
(b) GPT 4.1 Nano
Figure 15: LLaMA 3.1 8B Instruct and GPT 4.1 Nano Overrefusal Rates
K Directional Attack Success Rates Analysis (DirASR)
(a) LLaMA 3.1 8B Instruct
(b) GPT 4.1 Nano
Figure 16: LLaMA 3.1 8B Instruct and GPT 4.1 Nano vulnerability analysis using Directional ASR: ignore/completion, framing, and rhetorical strategy